Help! Point 3,-5 is on the terminal arm of angleX. The exact values of sinX and secX are...
Answer by stanbon(75887)

Point (3,-5) is on the terminal arm of angle theta.
---
If x = 3 and y = -5, r = sqrt[3^2+5^2] = sqrt[34]
------------------------------------

sin(theta) = y/r = 3/sqrt(34)
-----------------------------------
sec(theta) = r/x = sqrt(34)/(-5) = -sqrt(34)/5
===================================
Cheers,
Stan H.